OpenStudy (anonymous):

Solve the equation x-5/4= 3 for x. 17 7 12 10

OpenStudy (saifoo.khan):

x = 3 + 5/4

OpenStudy (anonymous):

??

OpenStudy (saifoo.khan):

??

OpenStudy (anonymous):

none

OpenStudy (anonymous):

they r all wrong

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

17-5=12/4=3